# generalised polynomial chaos

### v2

@@ III-1 Normal marginals
As $\lambda$ values are known exactly for normal distributions, the full process is first validated with arbitrary normally distributed marginals and linear bid functions. The correlation matrix $\bm{\Sigma}$ is fit from synthetic weather data of four different locations retrieved from renewables.ninja [9], however as there is no validation with real data, an arbitrary $\bm{\Sigma}$ could be used.
The polynomial basis is constructed from 9 monomials (all monomials of degree $\leq 1$). The quality of the gPC expansion is checked by drawing random samples from the Copula and evaluating both the bid functions and the expansion. The resulting distributions are displayed in Fig. 2. As constructed, the bid power distributions are indeed normally distributed. Moreover, the gPC expansion accurately approximates the bid functions. The maximum numerical error of the expansion was found to be on the order of $10^{-7}\%$
@@
Problem (24) is then solved with parameters $R^{(X)}$, $R^{(Y)}$, $T_{X\to Y}$, $T_{Y\to X}=100$ and $c_{1-8}=1$ to obtain optimal $\bm{x},\bm{y}$. The chance-constraints in (20) are validated by sampling both from the Copula and from the marginals in an uncorrelated. The resulting distributions are shown in Fig. 3. As expected, the optimization problem correctly satisfies the chance-constraints only for the correlated case. This results suggests, that the presented formalism works as intended.
Figure 3: Sampled procured power distributions computed with the gPC expansion in the case of normally distributed marginals. For comparison the PDFs both considering and not considering correlation were sampled. The empirical $1^{st}$ and $99^{th}$ percentiles are indicated for verification of the chance-constraints.
@@ III-2 Arbitrary marginals
Obviously, bid functions are in general not linear in the random variables and the marginals not normally distributed. For example, solar irradiance uncertainty is frequently modelled using Beta distributions due to its bounded support. For validation in a more realistic setting, the bid functions are replaced by simple polynomial models of PV and HP capability and the marginals by representative Beta distributions.
The polynomial basis was constructed from 21 physically relevant monomials. The computation of the basis and all coefficients was achieved in roughly 2 minutes total with a non-parallelised implementation and $k=15$ quadrature points per dimension. Validating the gPC expansions with sampling resulted in the bid power distributions shown in Fig. 4. The maximum expansion error was found to be on the order of $10^{-6}\%$
Problem (24) was solved with parameters $R^{(X)}$, $R^{(Y)}=1000$, $T_{X\to Y}$, $T_{Y\to X}=500$ and $c_{1-8}=1$. The constraints validation by sampling resulted in the distributions shown in Fig. 5. Due to the non-linearity of the bid functions and the non-normal, non-symmetric marginals, the constraints with $\lambda$ values for normal distributions are no longer met perfectly.
In principle, for each chance-constraint a local search for optimal $\lambda$ values is possible. Another option is the use of a distributionally robust $\lambda$, which however leads to conservative estimates as $\lambda_{1}^{rob},\lambda_{99}^{rob}=\mp\sqrt{\frac{1-0.01}{0.01}}=\mp 9.950$
@@
Figure 5: Sampled procured power distributions computed with the gPC expansion in the case of arbitrarily distributed marginals. For comparison the PDFs both considering and not considering correlation were sampled. The empirical $1^{st}$ and $99^{th}$ percentiles are indicated for verification of the chance-constraints.
@@ IV Conclusion
In this paper, we developed a tractable framework for modeling correlated stochastic energy resources using Polynomial Chaos Expansion combined with Gaussian Copulas. The approach enables to reformulate chance constraints on weighted sums of random functions as convex constraints. Several ideas were introduced to improve computational efficiency, including transformations for quadrature integration of Gaussian Copulas as well as precomputing expectation values for parallelization and effective dimensionality reduction.
The framework was validated on a numerical joint procurement problem with eight random variables and bids. First, using normally distributed marginals and linear bid functions, we showed that the reformulation correctly meets constraints for a correlated joint PDF. Then, for a more representative example with polynomial bid functions and Beta distributed marginals, a 21-dimensional orthonormal polynomial basis and corresponding PCE coefficients could be computed within two minutes. The optimization closely satisfies the chance constraints when using Gaussian quantile factors.
Key limitations include the reliance of dimensionality reduction techniques on the dimensionality of the random functions as well as the need for quantile factors $\lambda$ in the exact reformulation of chance constraints. While $\lambda$ is known for normal distributions and conservative robust bounds are available, its exact value remains generally unknown for arbitrary distributions.
Future work will focus on extending the framework to time-coupled problems, accounting for correlations across time steps. In addition, the formulation of joint chance constraints will be investigated.

### v3

@@ III-1 Normal marginals
As $\lambda$ values are known exactly for normal distributions, the full process is first validated with arbitrary normally distributed marginals and linear bid functions, together with an arbitrary correlation matrix $\bm{\Sigma}$. The polynomial basis is constructed from 9 monomials (all monomials of degree $\leq 1$). The quality of the gPC expansion is checked by drawing random samples from the Copula and evaluating both the bid functions and the expansion. The resulting distributions are displayed in Fig. 2. As constructed, the bid power distributions are indeed normally distributed. Moreover, the gPC expansion accurately approximates the bid functions. The maximum numerical error of the expansion is found to be on the order of $10^{-7}\%$, making the two distributions indistinguishable.
Problem (24) is solved with parameters $R^{(X)}$, $R^{(Y)}$, $T_{X\to Y}$, $T_{Y\to X}=100$ and $\gamma_{1-8}=1$ to get optimal $\bm{x},\bm{y}$. The chance-constraints in (20) are validated by sampling both from the Copula and from the marginal distributions assuming independence. The resulting distributions are shown in Fig. 3. As expected, the optimization problem tightly satisfies the chance-constraints only for the dependent case. This result shows that the presented formalism works as intended.
@@
Figure 3: Sampled procured power distributions computed with the gPC expansion in the case of normally distributed marginals. For comparison the PDFs both considering and not considering dependence are sampled. The empirical $1^{st}$ and $99^{th}$ percentiles are indicated for verification of the chance-constraints.
@@ III-2 Arbitrary marginals
In general, bid functions are not linear in the random variables and the marginals are not normally distributed. For example, solar irradiance uncertainty is frequently modelled using Beta distributions due to its bounded support. For validation in a more realistic setting, the bid functions are replaced by simple polynomials and the marginals by representative Beta distributions.
The polynomial basis is constructed from 21 relevant monomials. The computation of the basis and all coefficients is achieved in roughly 2 minutes total with a non-parallelised implementation and $k=15$ quadrature points per dimension. Validating the gPC expansions with sampling results in the bid power distributions shown in Fig. 4. The maximum expansion error is found to be on the order of $10^{-6}\%$, again with the two distributions indistinguishable.
Problem (24) is solved with parameters $R^{(X)}$, $R^{(Y)}=1000$, $T_{X\to Y}$, $T_{Y\to X}=500$ and $\gamma_{1-8}=1$. The validation of constraints satisfaction through sampling results in the distributions shown in Fig. 5. Due to the non-linearity of the bid functions and the non-normal, non-symmetric marginals, the constraints with $\lambda$ values for normal distributions are no longer met tightly. Note, that the overall computation time is dominated by the PCE process, while the choice of parameters has no noticeable impact, besides determining the feasibility of the problem.
@@
Figure 5: Sampled procured power distributions computed with the gPC expansion in the case of arbitrarily distributed marginals. For comparison the PDFs both considering and not considering dependence were sampled. The empirical $1^{st}$ and $99^{th}$ percentiles are indicated for verification of the chance-constraints.
@@ IV Conclusion
In this paper, we develop a tractable framework for modelling correlated random variables using PCE combined with Gaussian Copulas. The approach enables to reformulate chance constraints on weighted sums of random functions as convex constraints. Several ideas are introduced to improve computational efficiency, including transformations for quadrature integration of Gaussian Copulas as well as precomputing expectation values for parallelization and effective dimensionality reduction. The framework is validated on a numerical joint procurement problem with eight random variables and bids. First, using normally distributed marginals and linear bid functions, we show that the reformulation tightly meets constraints for a correlated joint PDF. Then, for a more representative example with polynomial bid functions and Beta distributed marginals, a 21-dimensional orthonormal polynomial basis and corresponding PCE coefficients could be computed within two minutes. The optimization closely satisfies the chance constraints when using Gaussian quantile factors. In addition to the methodological contributions, the proposed framework has relevance for system operators. For instance, TSOs engaged in interzonal AS provision or distribution system operators participating in reserve procurement can potentially benefit from the ability to model correlated stochastic resources in a tractable optimization setting. Key limitations include the reliance of dimensionality reduction techniques on the dimensionality of the random functions as well as the need for quantile factors $\lambda$ in the exact reformulation of chance constraints. While $\lambda$ is known for normal distributions and conservative robust bounds are available, its exact value remains generally unknown for arbitrary distributions. Future work will focus on extending the framework to time-coupled problems, accounting for correlations across time steps. In addition, the formulation of joint chance constraints will be investigated.